Grade Select GradeAlgebra

a circle passes through the points (-6,5),(-3,-4)and its radius is 5 units,find the equation of the circle

aradhya singh
9 Years agoGrade Select Grade

1 Answer

Vikas TU
9 Years ago
Since the circle passes through (-6,5),(-3,-4)
(-6-h)^2 + (5-k)^2 = (-3-h)^2 + (-4-k)^2
=> h+6 = 3k
=> h = 3k -6.
Now,
Since radius is 5 units and circle passes through (-3,-4)
(-3-h)^2 + (-4-k)^2 = 25
=> 9+(3k-6)^2 + 6(3k-6) + 16 + k^2 + 8k=25
10k^2 = 10k
K=0,1
If k =0
h=-6
Equation is (x+6)^2 + y^2 = 25
If k =1
h=-3
Equation is (x+3)^2 + (y-1)^2 = 25.